Paul Coddington

Northeast Parallel Architectures Center 111 College Place Syracuse University Syracuse, NY 13244, U.S.A

Phone: (315) 443-4883 FAX: (315) 443-1973 Email: paulc@npac.syr.edu

February 9, 1995

Selection Committee Ref R/141 Dept of Computer Science and Engineering University of Southampton

To whom it may concern,

I was in the same group as James Allwright at Southampton University, and knew something of his work although we did not collaborate on any projects. Dr Allwright has worked with me on a couple of projects at the Northeast Parallel Architectures Center at Syracuse University.

Dr Allwright is very bright and strong academically, although he is somewhat lacking in communication skills. I collaborated with him on a research project on parallel graph coloring algorithms and he did very good work. He has a good knowledge of parallel computing and parallel algorithms, strong programming skills, and very good technical writing abilities.

I believe he would perform very well on this project and I recommend him for this position.

Yours sincerely,

Paul Coddington